# A Simple Lifelong Learning Approach

### v19

@@ 6.4 Constant Resource Mode Operation
The binary distinction we made above, algorithms either build resources or reallocate them, is a false dichotomy, and biologically unnatural. In biological learning, systems develop from building to fixed resources, as they grow from juveniles to adults. To explore this continuum of amount of resources to grow, we experiment on FOOD1k 50X20 dataset using the constant resource mode operation of SiLLy-N as described in Section 3. We evaluate the performance of SiLLy-N for different number of encoder budget. As shown in Figure 9, performance of SiLLy-N saturates after $30$ encoders, though with only $5$ encoders, still demonstrates forward and backward transfer.
7 SiLLy-F on tabular data
@@
In this experiment, we experiment with SiLLy-F, an additional realization of our approach using random forests as encoders (described in Section 3). We flatten the CIFAR 10X10 data and use them as tabular data. We train two other best performing baseline algorithms, SiLLy-N and Model Zoo and use three fully connected hidden layers, each having $2000$ nodes, as encoders. As shown in Figure 10, SiLLy-F performs the best among all the approaches. This experiment shows our approach can be used as a general structure to do lifelong learning using other machine learning models as encoder.
8 Discussion
We introduced representation ensembling as a simple approach for lifelong learning. Two specific algorithms, SiLLy-N and SiLLy-F, achieve both forward and backward transfer, by leveraging resources learned for other tasks without undue computational burdens. In this paper, we have mainly focused on task-aware setting, because it is simpler. Future work will extend our approach to more challenging task-unaware settings. Our code, including code to reproduce the experiments in this manuscript, is available from http://proglearn.neurodata.io/.
Acknowledgements
The authors thank the support of the NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning (NSF grant 2031985). We also thank Raman Arora, Dinesh Jayaraman, Rene Vidal, Jeremias Sulam, Guillermo Sapiro, and Michael Powell for helpful discussions. This work is graciously supported by the Defense Advanced Research Projects Agency (DARPA) Lifelong Learning Machines program through contracts FA8650-18-2-7834 and HR0011-18-2-0025. Research was partially supported by funding from Microsoft Research and the Kavli Neuroscience Discovery Institute.
@@ Appendix A Evaluation Criteria
A.1 Representation Ensembling Algorithms
In this paper, we have proposed two representation ensembling algorithms, Simple Lifelong Learning Networks (SiLLy-N) and Simple Lifelong Learning Forests (SiLLy-F). The two algorithms differ in their details of how to update encoders and channels, but abstracting a level up they are both special cases of the same procedure. Let SiLLy-X refer to any possible representation algorithm. Algorithms  123, and  4 provide pseudocode for adding encoders, updating channels, and making predictions for any SiLLy-X algorithm. Whenever the learner gets access to a new task data, we use Algorithm 1 to train a new encoder for the corresponding task. We split the data into two portions — one set is used to learn the encoder and the other portion is called the held out or out-of-bag (OOB) data which is returned by Algorithm 1 to be used by Algorithm 2 to learn the channel for the corresponding task. Note that we push the OOB data through the in-task encoder and the whole dataset through the cross-task encoders to update the channel, i.e, learn the posteriors according to the new encoder. Then we use Algorithm 3 to replay the old task data through the new encoder and update their corresponding channels. Finally, while predicting for a test sample, we use Algorithm 4. Given the task identity, we use the corresponding channel to get the average estimated posterior and predict the class label as the $\operatornamewithlimits{argmax}$ of the estimated posteriors.
@@ A.1 Representation Ensembling Algorithms
4:(3) $\mathcal{D}_{n}^{t}=(\mathbf{x}^{t},\mathbf{y}^{t})\in\mathbb{R}^{n\times p}\times\{1,\ldots,K\}^{n}$triangleright training data for task $t$
@@
7:function SiLLy-X.update_channel($t,u_{t},\mathcal{D},\mathcal{V}$)
@@
9:         $v_{t^{\prime}}\leftarrow u_{t}$.update_channel($\mathbf{x}_{t^{\prime}},\mathbf{y}_{t^{\prime}},v_{t^{\prime}}$) $\triangleright$ update the old task channels
@@
12:     $\hat{y}=\operatornamewithlimits{argmax}(\hat{\mathbf{p}})$triangleright find the index of the elements in the vector $\hat{\mathbf{p}}$ with maximum value
@@
Table 1: Hyperparameters for SiLLy-N in CIFAR 10X10, Five Datasets, Split Mini-Imagenet, FOOD1k experiments. Note that we use the same hyperparameters for all the experiments.
@@
K (KNN channel) $\log_{2}(\text{number of samples per task})$
Table 2: Hyperparameters for SiLLy-F in tabular CIFAR 10X10 experiments.
@@
A.2 Reference Algorithm Implementation Details
@@
A.3 Training Time Complexity Analysis
@@
(15) $=\tilde{\mathcal{O}}(nT)$
A.4 Simulated Results
@@
A.5 Real Data Extended Results
FOOD1k and Mini-Imagenet datasets were obtained from https://www.kaggle.com/datasets/whitemoon/miniimagenet and https://github.com/pranshu28/TAG, respectively.
Table 3: Benchmark dataset details.
@@ A.5 Real Data Extended Results
Table 4: Hyperparameters for SiLLy-F in spoken digit experiment.
Hyperparameters Value
n_estimators ($275$ training samples per task) $10$
max_depth $30$
max_samples (OOB split) $0.67$
min_samples_leaf $1$
CIFAR 10x10 Repeated Classes
We also considered the setting where each task is defined by a random sampling of 10 out of 100 classes with replacement. This environment is designed to demonstrate the effect of tasks with shared subtasks, which is a common property of real world lifelong learning tasks. Supplementary Figure 2 shows transfer of SiLLy-F and SiLLy-N on Task 1.
@@
Figure 3: Extended results on the different vision experiments. This plot contains algorithms not shown in Figure 8.
Spoken Digit experiment
@@
Figure 5: Extended results on the Spoken Digit experiments. This plot contains algorithms not shown in Figure 8.

### v21

@@ VI-E Constant Resource Mode Operation
VI-E1 FOOD1k 50X20
The binary distinction we made above, algorithms either build resources or reallocate them, is a false dichotomy, and biologically unnatural. In biological learning, systems develop from building to fixed resources, as they grow from juveniles to adults. To explore this continuum of amount of resources to grow, we experiment on FOOD1k 50X20 dataset using the constant resource mode operation of SiLLy-N as described in Section IV. We evaluate the performance of SiLLy-N for different number of encoder budget. Performance of SiLLy-N saturates after $30$ encoders, though with only $5$ encoders, still demonstrates forward and backward transfer (Figure 8).
VI-E2 CORe50 110X5
To further evaluate the effectiveness of constant resource mode operation, we utilize a large-scale dataset: CORe50 [72], partitioned into $110$ tasks with $5$ classes each and $100$ training samples per class. To simulate a resource-constrained environment, we conducted all experiments on an Apple M1 Max chip with 64 GB of RAM. We set a budget of 20 encoders, as this configuration maintains high accuracy (above $90\%$) across all tasks. The fourth row of Figure 1 shows that SiLLy-N achieves higher $\mathsf{Transfer}$ operating in constant resource mode on a large dataset, that is, CORe50 compared to the smaller datasets discussed above. In particular, $\mathsf{BackwardTransfer}$ drops to zero after 20 tasks due to the encoder limit, leaving only $\mathsf{ForwardTransfer}$ for subsequent tasks. Moreover, we cannot run resource-growing algorithms like Model Zoo in this resource-constrained environment. The constant resource algorithm, such as LwF, completely forgets the first several tasks, and the corresponding task accuracies go as low as the chance level, $20\%$.
VI-F SiLLy-F on tabular data
@@
In this experiment, we experiment with SiLLy-F, an additional realization of our approach using random forests as encoders (described in Section IV). We flatten the CIFAR 10X10 data and use them as tabular data. We train two other best performing baseline algorithms, SiLLy-N and Model Zoo and use three fully connected hidden layers, each having $2000$ nodes, as encoders. As shown in Figure 9, SiLLy-F performs the best among all the approaches. This experiment shows our approach can be used as a general structure to do lifelong learning using other machine learning models as encoder.
VII Discussion
We introduced representation ensembling as a simple approach for lifelong learning. Two specific algorithms, SiLLy-N and SiLLy-F, achieve both forward and backward transfer, by leveraging resources learned for other tasks without undue computational burdens. Our work is well suited for federated learning scenarios, where each data center independently trains a model on its private data and shares only the encoders with other centers [73]. In this paper, we have mainly focused on task-aware setting, because it is simpler. Future work will extend our approach to more challenging task-unaware settings [74]. Our code, including the code for reproducing the experiments in this manuscript, is available from http://proglearn.neurodata.io/.
Acknowledgments
The authors thank the support of the NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning (NSF grant 2031985). We also thank Raman Arora, Dinesh Jayaraman, Rene Vidal, Jeremias Sulam, Guillermo Sapiro, Michael Powell, and Weiwei Yang for helpful discussions. This work is graciously supported by the Defense Advanced Research Projects Agency (DARPA) Lifelong Learning Machines program through contracts FA8650-18-2-7834 and HR0011-18-2-0025. Research was partially supported by funding from Microsoft Research and the Kavli Neuroscience Discovery Institute.
Co-first Author Contributions
J.D.: Conceptualization, Methodology, Software, Data Curation, Formal Analysis, Experimentation, Theoretical proof (see Appendix D), Writing &\&& Editing; J.T.V.: Conceptualization, Methodology, Supervision, Writing—Review &\&& Editing, and Funding Acquisition.
@@ Appendix A Evaluation Criteria
A-A Representation Ensembling Algorithms
A-A1 Model Architecture
In this paper, we proposed two representation ensembling algorithms, Simple Lifelong Learning Networks (SiLLy-N) and Simple Lifelong Learning Forests (SiLLy-F). The two algorithms differ in their details of how to update encoders and channels, but abstracting a level up they are both special cases of the same procedure. Let SiLLy-X refer to any possible representation algorithm. Algorithms  123, and  4 provide pseudocode for adding encoders, updating channels, and doing inference for any SiLLy-X algorithm.
A-A2 Data Preparation
Whenever the learner gets access to a new task data, we use Algorithm 1 to train a new encoder for the corresponding task. We split the data into two portions — in-task set and held out or out-of-bag set.
A-A3 Training Procedures
In-task set is used to learn the encoder and the indices of the out-of-bag (OOB) data which is returned by Algorithm 1 to be used by Algorithm 2 to learn the channel for the corresponding task. Note that we push the OOB data through the in-task encoder and the whole dataset through the cross-task encoders to update the channel, i.e, learn the posteriors according to the new encoder (see Algorithm 3). Finally, Algorithm 4 does inference on a new sample. Given the task identity, we use the corresponding channel to get the average estimated posterior and predict the class label as the $\operatorname*{arg\,max}$ of the estimated posteriors.
@@ A-A3 Training Procedures
4:(3) $\mathcal{D}_{n}^{t}=(\mathbf{x}_{t},\mathbf{y}_{t})\in\mathbb{R}^{n\times p}\times\{1,\ldots,K\}^{n}$triangleright training data for task $t$
@@
7:function SiLLy-X.update_old_channel($t,u_{t},\mathcal{D},\mathcal{V}$)
@@
9:         $v_{t^{\prime}}\leftarrow u_{t}$.update_channel($\mathcal{D}^{t^{\prime}},v_{t^{\prime}}$) $\triangleright$ update the old task channels
@@
12:     $\hat{y}=\operatorname*{arg\,max}(\hat{\mathbf{p}})$triangleright find the index of the elements in the vector $\hat{\mathbf{p}}$ with maximum value
@@
Figure 1: Error bars (interquartile range, i.e., IQR) for each dot in Figure 1 on different vision and speech benchmark datasets. Error bars for different performance statistics is negligible in comparison with the median performance shown in the main text Figure 1.
Figure 3: Performance of SiLLy-N on CIFAR 10X10 remains nearly unchanged for different number of decision trees per channel.
TABLE I: Hyperparameters for SiLLy-N in CIFAR 10X10, Five Datasets, Split Mini-Imagenet, FOOD1k experiments. Note that we use the same hyperparameters for all the experiments.
@@
n_estimators (decision forest channel) $20$
TABLE II: Hyperparameters for SiLLy-F in tabular CIFAR 10X10 experiments.
@@
A-B Reference Algorithm Implementation Details
@@
A-C Training Time Complexity Analysis
  1. 1.
    Each task has the same number of training samples.
  2. 2.
  3. 3.
    The number of epochs is fixed for each task.
  4. 4.
Assumption $3$ enables us to write time complexity as a function of the sample size. Table I summarizes the capacity, space and time complexity of several reference algorithms, as well as our SiLLy-N and SiLLy-F.
Lifelong learning methods are parametric if they have a representational capacity which is invariant to sample size and task number. Although the space complexity of some of these algorithms grow (because the number of the constraints stored by the algorithms grows, or they continue to store more data), their capacity is fixed. Thus, given a sufficiently large number of tasks with increasing complexity, in general, eventually all parametric methods will catastrophically forget. EWC [11], Online EWC [12], SI [14], and LwF [15] are all examples of parametric lifelong learning algorithms. Our fixed resource algorithms are also parametric.
Lifelong learning methods are semi-parametric if they have a representational capacity which grows slower than sample size. For example, if $T$ is increasing slower than $n$ (e.g., $T\propto\log n$), then algorithms whose capacity is proportional to $T$ are semi-parametric. ProgNN [21] is semi-parametric, nonetheless, its space complexity is $\tilde{\mathcal{O}}(T^{2})$ due to the lateral connections. Moreover, the time complexity for ProgNN also scales quadratically with $n$ when $n\propto T$. Thus, an algorithm that literally stores all the data it has ever seen, and retrains a fixed size network on all those data with the arrival of each new task, would have smaller space complexity and the same time complexity as ProgNN. DF-CNN [22] improves upon ProgNN by introducing a “knowledge base” with lateral connections to each new column, thereby avoiding all pairwise connections. Because these semi-parametric methods have a fixed representational capacity per task, they will either lack the representation capacity to perform well given sufficiently complex tasks, and/or will waste resources for very simple tasks.
Lifelong learning methods are non-parametric if they have a representational capacity which grow in proportion to the number of tasks or data samples. Table I shows the Indian Buffet Process for Weight Factors (IBP-WF) is a notable non-parametric approach alongside SiLLy-F.
Our proposed approaches, SiLLy-N and SiLLy-F, as we will discuss in details in Section IV, eliminate lateral connections between the columns of the network, thus reducing the complexity of the space to $\tilde{\mathcal{O}}(T)$. Moreover, our proposed approaches can adapt flexibly to any of the three categories based on the constraints of the application environment, as illustrated in Table I.
A-C1 Complexity analysis
@@
$=\tilde{\mathcal{O}}(nT)$ (13)
A-D Simulation Experiment Details
@@
This simulation setup facilitates the manipulation of task overlap, allowing for an examination of the transfer properties of our proposed approach under different levels of task similarity (see main text).
A-E Real Data Extended Experiments and Details
This section contains extended results on algorithms not shown in the main text (see Appendix Figure 5). FOOD1k and Mini-Imagenet datasets were obtained from https://www.kaggle.com/datasets/whitemoon/miniimagenet and https://github.com/pranshu28/TAG, respectively.
TABLE III: Benchmark dataset details.
@@ A-E Real Data Extended Experiments and Details
Split Mini-Imagenet
In this experiment, we have used the Mini-Imagenet dataset [65]. The dataset was split into $20$ tasks with $5$ classes each. Each task has $2400$ training samples and $600$ testing samples. As shown in Figure 5 right column, we get positive forward and backward transfer for SiLLy-N. However, although samples per task is lower compared to that of 5-dataset, it is still quite high. Hence, Model Zoo outperforms all the algorithms in this experiment.
5-dataset
In this experiment, we have used 5-dataset [65]. It consists of $5$ tasks from five different datasets: CIFAR-10 [67], MNIST, SVHN [76], notMNIST [77], Fashion-MNIST [78]. All the monochromatic images were converted to RGB format, and then resized to $3\times 32\times 32$. As shown in Appendix Table III, training samples per task in $5$-dataset is relatively higher than that of low data regime typically considered in lifelong learning setting. However, as shown in Figure 5 left column, SiLLy-N show less forgetting than most of the reference algorithms. On the other hand, Model Zoo shows comparatively better performance in relatively high task data size setup. Recall that SiLLy-N is based on bagging, and Model Zoo is based on boosting. It is well known that boosting often outperforms bagging when sample sizes are large  [^2] Authors in [79] shows that both bagging and boosting asymptotically converge to the Bayes optimal solution. However, for finite sample size and similar model complexity, we empirically find bagging approach to lifelong learning performs better than that of boosting when the training sample size is low whereas boosting performs better on large training sample size (See Figure 1). This is consistent with similar results in single task learning [80, 81, 41].
Overlapping Task Experiment
We considered the setting where each task is defined by a random sampling of 10 out of 100 classes with replacement in CIFAR 10x10. This environment is designed to demonstrate the effect of having overlapping tasks, which is a common property of real world lifelong learning tasks. Appendix Figure 4 shows positive transfer from other tasks to Task 1 for SiLLy-F and SiLLy-N.
@@
Figure 5: Performance of different lifelong learners on two vision datasets..
Spoken Digit Experiment Details
@@
Figure 7: Extended results on the Spoken Digit experiments. This plot contains algorithms not shown in main text Figure 1.